Instrument Log

 

Chill Mirror   C1                   Lamont, OK

 

Date

Technician

Comments

8/1/03

MR

This instrument needs to be removed from service.  Powered down the instrument at 1700 GMT.  Removed the instrument from service.

6/20/03

PAD

At 1545 GMT, flow rate was observed at 0.6.  Adjusted flow rate to 2.0 at 1548 GMT.  RTS at 1600 GMT.

5/12/03

JEM

The flow has dropped below 1.5 l/m.  Adjusted the flow to 2.0 l/m @1530 GMT.  RTS @1543 GMT.

4/28/03

JEM

The flow is only reading 1.4 and needs to be adjusted.  Adjusted the flow to 2.0 @1505 GMT.  RTS @1515 GMT.

4/10/03

JEM

The CMH flow rate is only 1.5.  The rate needs to be adjusted to 2.0.  Adjusted the flow rate to 2.0 @1547 GMT.  RTS @1600 GMT.

4/2/03

JEM

The flow has dropped to 1.5 l/m.  Adjusted the flow to 2.0 l/m @1617 GMT.  RTS @1630 GMT.

3/10/03

JEM

The flow has dropped to 1.3 l/m and needs to be adjusted.  Adjusted the flow from 1.3 l/m to 2.0 l/m @1516 GMT.  RTS @1530 GMT.

2/27/03

JEM

The flow has dropped to 1.4 l/m.  Adjusted the flow from 1.4 l/m to 2.0 l/m @1647 GMT.  RTS @1700 GMT.

2/21/03

MR

The instrument mentor has requested that the temp/hum probe on the chilled mirror be changed out.    R/R the *temp/hum probe.  RTS at 1703 GMT.  *Old - 55609 *New - R 53838

2/18/03

JEM

The flow has dropped to 1.4 l/m.  Adjusted the flow to 2.0 l/m @1447 GMT.  RTS @1500 GMT.

1/27/03

JEM

The flow needs to be increased because it only reads 1.2 l/m.  Adjusted the flow from 1.2 l/m to 2.0 l/m @1622 GMT.  RTS @1635 GMT.

1/13/03

JEM

The CMH flow meter is reading 1.5.  Adjusted the flow from 1.5 to 2.0 @1618 GMT.  RTS @1630 GMT.

12/16/02

JEM

The flow meter is only reading 1.4.  Adjusted the flow from 1.4 to 2.0 @1637 GMT.  RTS @1650 GMT.

10/25/02

JEM

The flow rate is only reading 1.5.  Adjusted flow from 1.5 to 2.0 @1615 GMT.  RTS @1630 GMT.

10/21/02

JEM

The flow rate is currently reading 1.0.  Adjusted the flow rate to 2.0 GMT.  RTS @1524 GMT.

10/11/02

JEM

The flow has dropped below 1.5 and is reading 1.3.  Adjusted flow to 2.0.  RTS @1445 GMT.

10/10/02

JEM

The flow has dropped below 1.5 and is reading 1.2.  Adjusted flow to 2.0.  RTS @1635 GMT.

8/2/02

CT

The flow meter needed to be adjusted. The flow meter was adjusted from 1.3 to the correct setting of 2.0 at 1509 GMT. RTS at 1524 GMT.

7/30/02

CT

The flow meter was registering at .6 on the gauge. The flow meter was adjusted from .6 to the correct setting of 2.0 at 1530 GMT. RTS at 1545 GMT.

7/19/02

CT

The Temp/Hum cap needed to be cleaned.  The Temp/Hum cap was cleaned at 1730 GMT. RTS at 1745 GMT. 

7/17/02

CT

The flow meter needed to be adjusted. The flow meter was adjusted from 1.4 to the correct setting of 1.9 at 1518 GMT. RTS at 1533 GMT.

7/15/02

CT

The flow meter needed to be adjusted. The flow meter was adjusted from .3 to the correct setting of 1.9 at 1411 GMT. RTS at 1426 GMT.

7/10/02

CT

The flow meter needed to be adjusted. The flow meter was adjusted from 1.3 to the correct setting of 1.9 at 1525 GMT. RTS at 1540 GMT.

7/8/02

CT

The flow meter was registering at 1.0 on the gauge. The flow meter was reset to the correct setting of 1.9 at 1440 GMT. RTS at 1455 GMT.

7/5/02

CT

The flow meter was registering at 1.4 on the gauge. The flow meter was adjusted to the correct setting of 1.9 at 1340 GMT. Status OK. RTS at 1355 GMT.

7/2/02

CT

The flow meter was registering at 1.3 on the gauge. The flow meter was adjusted to the correct setting of 1.9 at 1515 GMT. RTS at 1530 GMT.

6/28/02

CT

The flow meter needed to be adjusted. The flow meter was adjusted from .4 to 1.9 at 1757 GMT. RTS at 1813 GMT.

6/24/02

KT

The flow was reading 1.0 and needed to be adjusted.  At 1915 GMT adjusted the flow reading from 1.0 to 1.5.  RTS at 1930 GMT.  Status ok.

6/17/02

MR

Work order # 4327 requests that the temp/hum probe be changed out due to it producing readings in excess of 104%.  Changed out the *temp/hum probe.  RTS at 1930 GMT.  *Old - 24891/3   *New - 25550

6/17/02

CT

The flow meter needed to be adjusted. The flow meter was adjusted from .7 to the correct setting of 2.0 at 1606 GMT. RTS at 1621 GMT.

6/14/02

MR

Power restored.  System operational.

6/13/02

MR

The site has lost power due to storms in the area at 1415 GMT.  The instrument is shut down at this time.

6/10/02

CT

The flow meter needed to be adjusted. The flow meter was adjusted from 1.1 to 2.0 at 1520 GMT. RTS at 1535 GMT.

6/7/02

CT

The flow meter needed to be adjusted. The flow meter was adjusted from 1.1 to the correct setting of 2.0. Status OK. RTS at 1503 GMT.

6/3/02

CT

The flow meter was reading 1.0 on the gauge. The flow meter was adjusted to the correct setting of 2.0 at 1509 GMT. RTS at 1524 GMT.

5/28/02

CT

The flow meter was reading 1.2 on the gauge. The flow meter was adjusted from 1.2 to the correct setting of 2.0 at 1520 GMT. RTS at 1535 GMT.

5/23/02

CT

The flow meter was reading 1.0 on the gauge. The flow meter was adjusted from 1.0 to the correct setting of 2.0 on the gauge. Status OK. RTS at 1455 GMT.

5/13/02

CT

The flow meter was reading 1.1 on the gauge. Adjusted the flow meter to the correct setting of 2.0 at 1404 GMT. RTS at 1419 GMT.

5/6/02

CT

The flow meter was reading .6 on the gauge. Adjusted the flow meter to the correct setting of 2.0 at 1520 GMT. RTS at 1535 GMT.

4/22/02

CT

The flow meter was reading .8 on the gauge. Adjusted the flow meter from .8 to the correct setting of 2.0 at 1545 GMT. RTS at 1559 GMT.

4/12/02

CT

The flow meter was reading .5 on the gauge. Reset the flow meter to 2.0 on the gauge. RTS at 1517 GMT.

4/1/02

CT

The Flow meter was registering 0 on the gauge. Reset the Flow meter to the correct setting of 2.0 at 1622 GMT. RTS at 1637 GMT.

3/18/02

CT

The flow meter was reading 1.0. Adjusted the flow meter to the correct setting of 1.8 at 1855 GMT. RTS at 1910 GMT.

3/7/02

CT

The flow meter was reading 1.3 on the gauge. Reset the flow meter to the correct setting of 1.8 on the gauge at 1608 GMT. RTS at 1623 GMT.

3/4/02

CT

The flow meter was registering at .5 on the gauge.  Reset the flow meter to the correct setting of 1.8 at 1735 GMT. RTS 1750 GMT.

2/19/02

CT

The flow meter was reading 0.5. This is out of the targeted range. Reset the flow meter to 2.0 at 1926 GMT. RTS at 1941 GMT.

2/15/02

CT

The flow meter was reading 0.5. Reset the flow meter to 2.0 at 1651 GMT. RTS at 1705 GMT.

1/24/02

CT

The flow meter needed to be adjusted. The flow meter was adjusted on 1/24/02 at 1643 GMT. RTS at 1658 GMT.

1/23/02

CT

The flow meter was reading below the targeted 1.5 to 2.0 guidelines. Reset the flow meter to the correct setting at 1745 GMT. RTS at 1800 GMT.

1/22/02

CT

The flow meter was registering below the 1.5 to 2.0 guidelines established by the mentor. The flow meter was adjusted on 1/22/02 at 1705 GMT to the correct setting. RTS at 1720 GMT.

1/9/02

CJ

The Flow meter was reading below the 0.5 to 1.0 guidelines established by the mentor. Adjusted the flow meter to correct setting at 1709. Status OK. RTS at 1724 GMT.

1/8/02

CJ

The flow meter was reading below the 0.5 to 1.0 guidelines established by the mentor. Reset the flow meter to the correct setting at 1640 GMT. RTS at 1655 GMT.

12/26/2001

CJ

The temperature is not being displayed. T/S in progress. RTS at 1655 GMT.

12/24/2001

CJ

Flow meter gauge was reading below 0.5 to 1.0 guidelines. Adjusted flow meter to correct setting at 1545 GMT. RTS at 1600 GMT.

12/21/2001

CJ

Flow Meter needed to be adjusted. Adjusted Flow Meter to the specified range at 1925 GMT. RTS at 1940 GMT.

12/17/2001

CJ

Flow meter gauge was registering 0. Adjusted  the flow meter to the 0.5 to 1.0 range guideline that the mentor has established. RTS at 1915 GMT.

12/13/2001

CJ

The flow meter was out of the 0.5 to 1.0 range established by the mentor. Adjusted the flow meter to the correct setting at 1630 GMT. RTS at 1645 GMT.

12/12/01

MR

The instrument mentor contacted me about performing a service adjustment to the CMH.  This procedure is in the manual and is fairly simple.  I observed that an "S" was being displayed on the display indicating the unit needs a service adjustment as the manual states.  After reading the procedure, the CMH control box was opened and the instrument was placed in the heat mode for 1 min.  This mode will evaporate all remaining moisture on the mirror.  After 1 min I observed that there was 2 bars lit on the display.  Following the manual, I turned the adjustment screw clockwise until only one bar was lit.  I then placed the balance switch to INIT position to reset the instrument.  After the system reset itself I noticed the "S" indicator had gone out as expected.  RTS at 2100 GMT.

12/11/01

CJ

Flow Meter was registering out of the 0.5 to 1.0 guidelines that have been established by the mentor. Reset flow meter @ 1825 GMT. RTS @ 1840 GMT.

6/27/01

MR

The mentor indicated that the instrument was not working properly and wanted me to check the air flow.  The air flow meter showed 0 scfm.  Turned the knob back and forth and the air flow returned.  Adjusted the air flow to .8 scfm.  RTS at 1315 GMT.

12/11/00

CAM

Due to a hard drive failure on the SDS Development system, data stopped being collected from the Chilled Mirror at 2131 GMT on 12/6/00. The hard drive was replaced and normal data collections resumed at 2033 GMT on 12/11/00 after all the necessary configurations were set.

12/11/00

MR

Cleaned the mirror as per mentors instruction at 2015 GMT.

12/07/00

KT

At 1745 GMT performed cleaning of the chilled mirror. RTS at 1753 GMT.

8/16/00

MR

SDS operator notified me that the CM was not collecting since yesterdays power cycle. Power cycled the CM at 1515 GMT. Observed on the clouds page that the instrument had recoved at 1730 GMT.

8/15/00

MR

Cleaned the mirror. The mentor also requested that the instrument be power cycled. Power cycled the instrument at 2030.

2/15/2000

MR

Reloaded the program into the data logger as per mentors request at 2005.

1/25/2000

MR

Instrument mentor has sent a new program to be downloaded into the data logger. Downloaded the program at 1700 GMT. Status O.K.

12/6/99

MR

The instrument mentor wants the mirror cleaned and 2 sets of readings taken compared to a hand held temp/hum probe. Took the first set of readings between 1648-1652 with the handheld. Cleaned the mirror and waited till 1750 to take second set of readings. Recorded the readings and sent them to the mentor.

8/24/99

MW

The instrument mentor requested that the chilled mirror be cleaned. Powered the chilled mirror off at 2017 GMT and cleaned the mirror. Powered back on at 2020 GMT.

8/9/99

MW

The instrument mentor requested that the Chilled mirror be powered off because it was not operating correctly. No further information given. The mentor indicated that he would be here at the CF later this week to correct the problem. Unplugged the grey power cord from the power strip as per mentor instruction at 1830 GMT.

7/12/99

MW

The instrument mentor requested that the CMH air intake flow rate be checked, the mirror be cleaned, and a plastic funnel be attached to the air intake hose to prevent water from entering the system through the intake hose. The flow rate was at 1.5 on the scale. Adjusted the flow to 1.0 at 1326 GMT as per mentor instructions. Cleaned the mirror at 1327 GMT. Do not have a funnel at this time, so that procedure has not been completed yet. Will install the funnel when one becomes available.

6/11/99

MW

The instrument mentor requested that the chilled mirror be cleaned. Opened the hygrometer and cleaned the chilled mirror with a cotton swab and the supplied blue liquid 1530-1545 GMT.

4/21/99

CAM

Instrument mentor was on site today for system modifications. Instrument was RTS at 1800 GMT but has not collected successfully following the maintenance. T/S found that the adapter which is between and supplies power to the bit driver and SC32A was reversed. As a result the bit driver and SC32A were not being powered on the correct pins. Turned the adapter around and RTS at 2030 GMT. The following collection was successful at 2031 GMT.

4/21/99

SR (Mentor)

Scott Richardson on site to replace chilled mirror system. Also removed the spare *T/RH sensors that were installed on 2/22/99. *Old - 53836 *Old 53837. New CR10 SN 26105. New SM716 SN 5892. T/RH sensor 53838.

4/16/99

CAM

SDS comlog reports indicated that the chilled mirror had not successfully collected since 0131 GMT. The error message 'Could not get attention of logger' was generated. Monitored the chilled mirror data logger using a laptop computer at 1420 GMT. Nearly all channels were reporting 00000. Channel 1 was reading -99999 and channel 45 was reading 99999. Channel 46 appeared to be reporting good values and channel 47 (battery) was reporting 11.8. Contacted Scott Richardson who advised the program problem first occurred following a power outage several days ago. He was aware of the program problem and plans to come and next week and replace the entire system with a new one. However, the system had continued to collect data files up until this morning. Reconnected all data lines at 1435 GMT. The following collection was successful at 1531 GMT. Suspect a loose connection was the cause of the missed collections and Scott will be out next week to correct the program problem. RTS at 1435.

2/22/99

SR (Mentor)

Added 2 new Rotronics HP043 *T/RH sensors. The sensors will ultimately be used with the 2 additional Chilled Mirror systems. Changed data logger program to ingest additional sensor data. Did not clean mirror, service not required. *New - 53836 *New - 53837

12/2/98

CM

The instrument mentor requested that the filter be removed from the chilled mirror air line. Disconnected the air line from the filter and left the line open to outside air. Secured the line below the box in such a manner as to prevent water from entering.

11/16/98

MW

The instrument mentor notified site operations on 11/16/98 that no data files were being collected from the BBSS CMH. Tipped the instrument from Dev10 and began receiving garbled characters back. Opened the instrument enclosure and removed the SC32A at 1740 GMT. Opened the SC32A and found the internal jumper was in the open position. Closed the internal jumper and placed the SC32A back into service. Checked communications from Dev10 at 1755 GMT and was able to receive stars back from the data logger. The SDS began receiving data files on the next collection at 1832 GMT.

6/16/98

CAM

SDS trouble report indicated no files had been received from the BBSS CMH since 6/14/98 at 1036 GMT. SDS comlog indicates 'Could Not Get Attention of Logger'. Found the receive light on constantly at the instrument end. Power cycled bit driver but receive light again lit constantly. R/R bit driver with a spare at 1315 GMT on 6/16/98. The spare bit driver receive light did not stay lit. The following collection failed with same error at 1335 GMT. Collected all data from the storage module using a laptop and PC208W software at 1358 GMT. Erased storage module and restored program at 1410 GMT. The following two collections also failed at 1435 and 1535 GMT. Removed the bit driver that had been installed previously and replaced with the original at 1600 GMT. The following collection was successful at 1635 GMT. It appears that erasing the storage module and restoring the program resolved the problem as the original bit driver is back in place.

 6/3/98

CAM

SDS TR indicated no files had been received from the Chilled Mirror since 0035 GMT on 6/2 with the message 'Could not get data from logger'. Monitored instrument and everything appeared O.K. Checked status of storage module using PC208W. Status check indicated that the module was full. Collected all data and then erased module and restored program at 2025 GMT. The following collection at 2035 failed with the same error message. Found that the bit driver transmit light was constantly flashing. R/R *bit driver at 2100 GMT. The following collection was successful at 2135 GMT. RTS status O.K. *Old - 34553 *New - 35935

4/22/98

MR

Work order # 2789 states that the BBSS Chilled Mirror has not sent any files to the SDS. Downloaded the program Gee41.dld into the storage module at 1936. This did not fix the problem. The storage module was observed as full. Collected all data from the storage module using the program PC208W then erased the storage module using the same program. SDS started to collect data at 0625 on 4/23/98.